OpenStudy (anonymous):

If f(x)=In(e/x^n), and n is a constant, what's f'(x)? I think it's x^n/e, am I right?

OpenStudy (anonymous):

=lne -lnx^n (log rules) =1-n*(lnx) now differentiate =0-n/x